y varies directly as the square of R. If y = 196 when R = 7 find y when R - 3
Answer by Alan3354(30983)

y varies directly as the square of R. If y = 196 when R = 7 find y when R - 3
---------------
I suppose you mean when R = 3.
y = kR^2
196 = k*7^2 = 49k
k = 4
------
@R = 3, y = 4*3^2 = 36